OpenStudy (anonymous):

which of the following ratio are eequivalent 6:9 choose all that apply. 3:2 ,18/27, 24/32, 30:45. I GREEN COME HELP ME PLEASE

OpenStudy (igreen):

@noterdame13 Tag me using the @ sign, then I'll come.

OpenStudy (anonymous):

WHAT SIGHN

OpenStudy (igreen):

The @ sign, then type my name in. like this: @iGreen

OpenStudy (igreen):

Press Shift and 2 to use it.

OpenStudy (igreen):

Anyway. 6 * 3 = 18 9 * 3 = 27 So 18/27 is an answer. 6 * 4 = 24 What's 9 * 4? @noterdame13

OpenStudy (anonymous):

36

OpenStudy (igreen):

Yep, so 24/36 is also an answer. So our answers are: 18/27 and 19/36

OpenStudy (igreen):

But there isn't a 19/36, only a 19/32, which is WRONG. 6 * 5 = 30 9 * 5 = ? @noterdame13

OpenStudy (anonymous):

40

OpenStudy (igreen):

No, check again.

OpenStudy (anonymous):

19/36

OpenStudy (igreen):

9 * 5 = ?

OpenStudy (anonymous):

45

OpenStudy (igreen):

Yep, so that gives us 30/45, which is also an answer. So our answers are 18/27 and 30/45.